OpenStudy (anonymous):

Please Help! Solve the equation: -14 = -(-2x + 2)

OpenStudy (anonymous):

Please show step. :)

OpenStudy (anonymous):

steps*

OpenStudy (ivettef365):

first solve parenthesis as follows: -14 = -(-2x + 2) -14 = +2x -2

OpenStudy (ivettef365):

now add 2 on both sides, what do you get ?

OpenStudy (anonymous):

-12?

OpenStudy (ivettef365):

ok, so you have -12 = 2x now divide by 2

OpenStudy (anonymous):

Would it be x = -6?

OpenStudy (phi):

2x= -12 means 2 x's add up to -12. 1 x will be -6 to solve using algebra rules, you divide both sides by 2 \[ \frac{2x}{2} = \frac{-12}{2} \] on the left side 2/2 is 1 so it simplifies to 1 * x or just 1 on the right side you get -6 x = -6

OpenStudy (ivettef365):

correct

OpenStudy (phi):

on the left side 2/2 is 1 so it simplifies to 1 * x or just x on the right side you get -6 x = -6

OpenStudy (anonymous):

Okay, thank you so much guys! I really appreciate the help! :)